# Strong parity-violation effects induced by large-amplitude motions: A quantum-dynamics study of substituted chiral methanols

### v1

@@ Abstract
Here, the PV shift in the torsional motiosns of the CXYZ unit in CXYZOH (X, Y, Z = H, F, Cl, Br, I) molecules was investigated. Strong PV enhancements are observed due to the localization of the torsional wavefunction to the geometries exhibiting opposite signs. The results obtained for CHBrIOH, CFBrIOH, CFClBrOH, and CClBrIOH molecules are reported as exxamples, and those for the other molecular species are reported in the Supporting Information (SI).
Figure 1 shows the overflow for obtaining the PV frequency of the torsional states.
@@
The minimum energy on the PEC was obtained using the simplex code [37], and PEC was shifted so that its bottom could become zero .
@@
The Cartesian coordinates optimized in Step 1 with respect to were transformed into the relaxed reference internal coordinates. First, we extracted and from the Cartesian coordinate defined in eq 4 and obtain the torsional angle (; eq 20). Further, (, which covers ) is obtained from , and these discrete coordinates are interpolated to obtain . Additionally, the horizontal axis of the potential energy is also transformed (, while is interpolated.
@@
The following definition was used to obtain the body-fixed Cartesian coordinate (Figure 2), following ref 37
where the atomic number increases from atoms X-Z. The frame for the internal rotation of the CXYZ unit was also defined based on the body-fixed Cartesian coordinate definition (eq 4). Furthermore, the center of mass was fixed at the origin of the body-fixed coordinate system, which offers an additional constant shift vector to each vector in eq 4. These coordinate definitions were used in the vibrational calculations to prevent the rotation of the frame.
@@
This definition facilitated the identification of as the single LAM in the system, with the small-amplitude motions (SAMs). This extraction technique for was initially proposed in refs [38, 39]. For subsequent convenience, we define the following compact notation for the internal coordinates,
@@
where is the unit vector in the body-fixed Cartesian coordinates. In the actual calculations, the -vectors are obtained at the discrete-variable representation (DVR) grid points (), i.e., the matrix elements are constructed based on rather than in eqs 24 and 25. The reference internal coordinates for obtaining are relaxed at each , where is the geometry that minimizes at a given , . Employing the relaxed coordinates indicates that the moment of inertia in this study depends on . The wavefunctions and energies associated with eq. 22 were obtained in a variational manner.
@@
Here, is the Fermi coupling constant, is the weak charge of the nucleus , with its neutron number and nuclear charge . The Weinberg mixing angle [49] was also employed. is the Dirac matrix that results in the coupling between the large and small components of the wavefunction. Additionally, is the charge density of nucleus for which the Gaussian-type model [50] is employed in this study. As the PV interaction is mediated by the heavy boson, the contact interaction between and an electron offers a good approximation. Although Eq. 26 is a four-component expression, it is numerically transformed into a two-component framework (picture-change [51]) in this study. As the sign of the becomes opposite after parity inversion, the energy difference due to the PV interaction between the - and -configuration can be expressed by
@@
where expresses the value of at the nuclear configuration, . We obtained the expectated value with respect to the DVR vibrational wavefunction. This indicates that the error due to the truncation of the potential (e.g., the harmonic oscillator and anharmonic correction), which is traditionally employed in this field, is not observed in this manner.
@@
The vibrational Schrödinger equation in a reduced-dimension form was solved using the GENIUSH program [41, 42], which was recently reviewed in ref 52. The derivative in eq 24 was obtained numerically (two-point central derivative) using the step size bohr. Furthermore, 37 Fourier-DVR grid points, which is a sufficient number of grid points, were used to solve the vibrational Schrödinger equation, as presented in Table S1 of the SI. The TENSINT program [53, 54] was used to obtain at each torsional energy level.
@@
Geometry optimization was carried out using the Gaussian16 software package [55]. For the Br and I atoms, we employed the psuodopotentials ECP28MDF and ECP10MDF, respectively, and the cc-pVTZ-PP basis sets [56, 57]. For the H, C, O, F, and Cl atoms, we employed the Def2TZVPP basis sets [58]. The PBE1PBE functional [59, 60] was used for the density functional theory (DFT) functional. Dispersion effects were considered within the semi-empirical D3 van der Waals corrections [61].
@@
To calculate the total electronic energy and , we employed a development version (git hash: 3970bcf) of the DIRAC program package [62, 63]. We constructed a two-component relativistic Hamiltonian using the exact two-component (X2C) Hamiltonian [64] for the one-electron operators and the extended atomic mean-field (eAMF-X2C) approach [65] to consider the picture-change effect of the two-electron Coulomb and Gaunt terms. The PBE0 functionals [59, 60] were employed for the electronic structure calculations, and the dyall.3zp basis sets [66, 67, 68] were employed for all the elements in the molecules in the uncontracted form. Several tight exponents, which were optimized using the simplex method [69], were also added (Table S2 of the FIGSupporting Information). The Gauss-type finite nuclear-charge model [50] was employed for the relativistic calculations.
The dataset associated with this manuscript (the outputs from calculations and scripts to obtain the figures) is provided in ref 70.
@@
Figure 3 shows the criticality of the geometry-relaxation effect in the generation of the 1D PEC and curve. The fixed-geometry model cannot reproduce the results of the relaxed structure, as some wells in the PEC do not appear in the fixed-geometry model. Although the situation is better in , the fixed-geometry model overestimates the PV sensitivities in many regions. From a different perspective, geometries deviating from the optimized torsional path can display larger PV sensitivities, and it encourages full-dimensional quantum dynamics calculations in future studies. The relaxed structure model was used for the following discussion. Although the shape of the curve appeared complicated, two sinusoidal curves (a local minimum and a local maximum at 0∘-180∘ and 180∘-360∘, respectively) can be found for some target molecules (see Figures 4 and S1-S6 in the SI).
@@
The enhancement of the PV shift induced by the torsional motion can be visually demonstrated. Figure 4 shows the PEC, curve, and torsional wavefunctions of some CXYZOH molecules. The curves take various values along , and the torsional wavefunctions located in different wells can exhibit different values. This yields large values in the transition across the two wells. This is in contrast to SAMs, where the small displacement from the equilibrium structure contributes to the vibrational wavefunction, and the two states exhibit a similar magnitude. Following ref 29, the contribution from the term is small, yielding a linear-like curve with respect to the vibrational coordinate (). The value and vibrational band origin for each torsional state are presented in Table 1. Assuming the state control was possible, the maximum sensitivity of 3.1 Hz would be achieved in CHBrIOH. The PV frequencies from the ground state exceeded more than 1.0 Hz in the highlighted transitions of the CHBrIOH and CFBrIOH molecules. Compared with the CHFClBr molecule with = 1.867 (–3.781) mHz for the C-F stretching (FCCl bending) fundamental modes [71], approximately 100- (50-)fold enhancement was observed in CFClBrOH (186.4 mHz) without including a heavy element. The necessary precisions (the ratio between the PV shift and torsional frequency) for detecting the PV shift are and for CHBrIOH and CFClBrOH, respectively. The reliability of the enhancement mechanism in these molecules is detailed in section S4 of the SI.
@@
Some torsional transitions appear in the terahertz (THz) range, where lasers are actively developed to achieve stable frequencies [72], as in the case of quantum cascade lasers [73, 74]. However, the transition to the SAM-LAM coupling mode can reach the wavelengths in the infrared range, where CO2 laser can reach. In the case of the H-containing species, PV-sensitive transitions are observed in the infrared range, such as the transition between #1 and #5 of CHBrIOH presented in Table 1. To the best of our knowledge, these substituted methanol molecules have not been synthesized; however, they may be formulated in the gas phase via laser ablation, which is widely employed (e.g., refs 75, 76, 77, 78, 79).
@@ Acknowledgements.
The author thanks financial support from the European Union’s Horizon 2022 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101105452 (QDMAP) and the JSPS KAKENHI (grant no 21K14643). The author used the computer resource offered under the category of General Projects by the Research Institute for Information Technology, Kyushu University, and the FUJITSU Supercomputer PRIMEHPC FX1000 and FUJITSU Server PRIMERGY GX2570 (Wisteria/BDEC-01) at the Information Technology Center, The University of Tokyo. The author thanks Edit Mátyus (Budapest) for valuable discussions and Alberto Martń Santa Daría (Salamanca) for the support in using the TENSINT code. The author thanks Editage (www.editage.com) for English language editing.References
@@
- Aucar et al. [2022] I. A. Aucar, Y. Chamorro, and A. Borschevsky, Parity-violating contributions to nuclear spin-rotation interactions and to NMR shielding constants in tetrahedral molecules, Phys. Rev. A 106, 10.1103/physreva.106.062802 (2022).
- [49] 0.2319 is the default value of the version of the DIRAC code employed in this study. The value of depends on the renormarization prescription, see ref Zyla2020PTEP_reiview.
@@
- [63] DIRAC, a relativistic ab initio electronic structure program, Release DIRAC22 (2022), written by H. J. Aa. Jensen, R. Bast, A. S. P. Gomes, T. Saue and L. Visscher et al. (available at http://dx.doi.org/10.5281/zenodo.6010450, see also http://www.diracprogram.org), accessed Sep. 29, 2022.
@@
- Sunaga [2024] A. Sunaga, Strong parity-violation effects induced by large-amplitude motions: A quantum-dynamics study of substituted chiral methanols: Dataset, Zenodo https://doi.org/10.5281/zenodo.14030414 (2024).
@@
- Augenbraun et al. [2020] B. L. Augenbraun, J. M. Doyle, T. Zelevinsky, and I. Kozyryev, Molecular asymmetry and optical cycling: Laser cooling asymmetric top molecules, Physical Review X 10, 31022 (2020), 2001.11020 .
- Persinger et al. [2023] T. D. Persinger, J. Han, A. T. Le, T. C. Steimle, and M. C. Heaven, Electronic spectroscopy and ionization potentials for YbOH and YbOCH3, Phys. Rev. A 107, 10.1103/physreva.107.032810 (2023).
@@ I Employed masses
II Convergence with respect to the DVR grid points
@@ III Tight exponents
IV Reliability of the enhancement mechanism
The PV enhancements reported in the main text would not be due to the approximation employed in this study. The qualitative accuracy of the reduced dimensional models that consider only large-amplitude motion(s) has been confirmed in previous studies [81, 82, 83, 41, 84]. The PBE0 functional offers an acceptable quality (maximum absolute error of 0.5 kcal/mol from the CCSD(T) level) for the torsional potential energy including halogens (Figure 5 of ref [85]). The torsional coordinates that offer the maximum, zero, and minimum values do not significantly depend on the methodology choice, e.g., the sinusoidal curves of H2X2 (Fig. 1 of [20]), and the torsional curves of dichlorinedioxide and substituted allenes [86]. Although the maximum and minimum values depend on the methodology, the error due to PBE0 would be much smaller than the enhancement, as the maximum difference between at the PBE0 and CCSD levels of theory is approximately 11% for the H2X2 (X = O, S, Se, Te, Po) molecules [69].
V PECs, wavefunctions, and PV curves
@@
Figures 6-10 display the potential energy curves (PECs), 1D torsional wavefunctions, and parity-violating energy () curves of the molecules that are not shown in the main text. In these figures, the blue and green solid lines indicate the odd (ground, 3rd, …) and even-numbered (2nd, 4th, …) torsional state levels. The horizontal axis represents the torsional angle () defined in eq 2 of the main text. An example of the direction of the internal rotation of the CXYZ unit is shown in Figure 3 of the main text. The atomic number increases from atom X to Z, and the wells correspond to , respectively, corresponding to (H-O-C-Y), (H-O-C-Z), and (H-O-C-X), respectively.
Tables 6-10 present the vibrational energies and PV frequencies of the target molecules. In these tables, ( = X, Y, Z = H, F, Cl, Br, I) represents the quantum number of the well where the dihedral angle -C-O-H is close to at the corresponding (approximate) values (in degrees). “-” refers to mixed states that are delocalized into the three wells.

### v2

@@ I Introduction
Here, the PV shift in the torsional motions of the CXYZ unit in CXYZOH (X, Y, Z = H, F, Cl, Br, I) molecules was investigated. Strong PV enhancements are observed due to the localization of the torsional wavefunction to the geometries exhibiting opposite signs. The results obtained for CHBrIOH, CFBrIOH, CFClBrOH, and CClBrIOH molecules are reported as examples, and those for the other molecular species are reported in the Supplementary Material (SM).
II Theory
Figure 1 shows the flow diagram for obtaining the PV frequency of the torsional states.
@@ II Theory
The minimum energy on the PEC was obtained using the simplex code [37], and the PEC was shifted so that its bottom could become zero .
@@
The Cartesian coordinates optimized in Step 1 with respect to were transformed into the relaxed reference internal coordinates. First, we extracted and from the Cartesian coordinate defined in Eq. 4 and obtain the torsional angle (; Eq. 20). Further, (, which covers ) is obtained from , and these discrete coordinates are interpolated to obtain . Additionally, the horizontal axis of the potential energy is also transformed (, while is interpolated.
@@
The following definition was used to obtain the body-fixed Cartesian coordinate (Figure 2), following Ref. 37
where the atomic number increases from atoms X-Z. The frame for the internal rotation of the CXYZ unit was also defined based on the body-fixed Cartesian coordinate definition (Eq. 4). Furthermore, the center of mass was fixed at the origin of the body-fixed coordinate system, which offers an additional constant shift vector to each vector in Eq. 4. These coordinate definitions were used in the vibrational calculations to prevent the rotation of the frame.
@@
This definition facilitated the identification of as the single LAM in the system, with the small-amplitude motions (SAMs). This extraction technique for was initially proposed in Refs. [38, 39]. For subsequent convenience, we define the following compact notation for the internal coordinates,
@@ II.1 Vibrational Hamiltonian
where is the unit vector in the body-fixed Cartesian coordinates. In the actual calculations, the -vectors are obtained at the discrete-variable representation (DVR) grid points (), i.e., the matrix elements are constructed based on rather than in eqs 24 and 25. The reference internal coordinates for obtaining are relaxed at each , where is the geometry that minimizes at a given , . Employing the relaxed coordinates indicates that the moment of inertia in this study depends on . The wavefunctions and energies associated with Eq. 22 are obtained in a variational manner.
II.2 Parity-violating Hamiltonian
@@
Here, is the Fermi coupling constant, is the weak charge of the nucleus , with its neutron number and nuclear charge . The Weinberg mixing angle [49, 50] was also employed. is the Dirac matrix that results in the coupling between the large and small components of the wavefunction. Additionally, is the charge density of nucleus for which the Gaussian-type model [51] is employed in this study. As the PV interaction is mediated by the heavy boson, the contact interaction between and an electron offers a good approximation. Although Eq. 26 is a four-component expression, it is numerically transformed into a two-component framework (picture-change [52]) in this study. As the sign of the becomes opposite after parity inversion, the energy difference due to the PV interaction between the - and -configuration can be expressed by
@@
where expresses the value of at the nuclear configuration, . We obtained the expected value with respect to the DVR vibrational wavefunction. This indicates that the error due to the truncation of the potential (e.g., the harmonic oscillator and anharmonic correction), which is traditionally employed in this field, is not observed in this manner.
III Computational details
@@
The vibrational Schrödinger equation in a reduced-dimension form was solved using the GENIUSH program [41, 42], which was recently reviewed in Ref. 53. The derivative in Eq. 24 was obtained numerically (two-point central derivative) using the step size bohr. Furthermore, 37 Fourier-DVR grid points, which is a sufficient number of grid points as presented in Table S1 of the SM, were used to solve the vibrational Schrödinger equation. The TENSINT program [54, 55] was used to obtain at each torsional energy level.
@@
Geometry optimization was carried out using the Gaussian16 software package [56]. For the Br and I atoms, we employed the pseudopotentials ECP10MDF and ECP28MDF, respectively, and the cc-pVTZ-PP basis sets [57, 58]. For the H, C, O, F, and Cl atoms, we employed the Def2TZVPP basis sets [59]. The PBE1PBE functional [60, 61] was used for the density functional theory (DFT) functional. Dispersion effects were considered within the semi-empirical D3 van der Waals corrections [62].
@@
To calculate the total electronic energy and , we employed a development version (git hash: 3970bcf) of the DIRAC program package [63, 64]. We constructed a two-component relativistic Hamiltonian using the exact two-component (X2C) Hamiltonian [65] for the one-electron operators and the extended atomic mean-field (eAMF-X2C) approach [66] to consider the picture-change effect of the two-electron Coulomb and Gaunt terms. The PBE0 functionals [60, 61] were employed for the electronic structure calculations, and the dyall.3zp basis sets [67, 68, 69] were employed for all the elements in the molecules in the uncontracted form. Several tight exponents, which were optimized using the simplex method [70], were also added (Table S2 of the SM). The Gauss-type finite nuclear-charge model [51] was employed for the relativistic calculations.
IV Results and discussion
@@
Figure 3 shows the criticality of the geometry-relaxation effect in the generation of the 1D PEC and curve. The fixed-geometry model cannot reproduce the results of the relaxed structure, as some wells in the PEC do not appear in the fixed-geometry model. Although the situation is better in , the fixed-geometry model overestimates the PV sensitivities in many regions. From a different perspective, geometries deviating from the optimized torsional path can display larger PV sensitivities, and it encourages full-dimensional quantum dynamics calculations in future studies. The relaxed structure model was used for the following discussion. Although the shape of the curve appeared complicated, two sinusoidal curves (a local minimum and a local maximum at 0∘-180∘ and 180∘-360∘, respectively) can be found for some target molecules (see Figures 4 and S1-S6 in the SM).
@@
The enhancement of the PV frequency shift induced by the torsional motion can be visually demonstrated. Figure 4 shows the PEC, curve, and torsional wavefunctions of some CXYZOH molecules. The curves take various values along , and the torsional wavefunctions located in different wells can exhibit different values. This yields large values in the transition across the two wells. This is in contrast to SAMs, where the small displacement from the equilibrium structure contributes to the vibrational wavefunction, and the two states exhibit a similar magnitude. Following Ref. 29, the contribution from the term is small, yielding a linear-like curve with respect to the vibrational coordinate (). The value and vibrational band origin for each torsional state are presented in Table 1. Assuming the state control was possible, the maximum sensitivity of 3.1 Hz would be achieved in CHBrIOH. The PV frequencies from the ground state exceeded more than 1.0 Hz in the highlighted transitions of the CHBrIOH and CFBrIOH molecules. Compared with the CHFClBr molecule with = 1.867 (–3.781) mHz for the C-F stretching (FCCl bending) fundamental modes [71], approximately 100- (50-)fold enhancement was observed in CFClBrOH (186.4 mHz) without including a heavy element. The necessary precisions (the ratio between the PV shift and torsional frequency) for detecting the PV shift are and for CHBrIOH and CFClBrOH, respectively.
These PV enhancements would not be due to the approximation employed in this study. The qualitative accuracy of the reduced dimensional models that consider only large-amplitude motion(s) has been confirmed in previous studies [72, 73, 74, 41, 75]. The PBE0 functional offers an acceptable quality (maximum absolute error of 0.5 kcal/mol from the CCSD(T) level) for the torsional potential energy including halogens (Figure 5 of ref [76]). The torsional coordinates that offer the maximum, zero, and minimum values do not significantly depend on the methodology choice, e.g., the sinusoidal curves of H2X2 (Fig. 1 of Ref. 20), and the torsional curves of dichlorinedioxide and substituted allenes [77]. Although the maximum and minimum values depend on the methodology, the error due to PBE0 would be much smaller than the enhancement, as the maximum difference between at the PBE0 and CCSD levels of theory is approximately 11% for the H2X2 (X = O, S, Se, Te, Po) molecules [70]. The PV enhancement due to the torsional motion was also observed with the CAM-B3LYP* functional [78], as shown in Sec S6 of the SM.
@@
Some torsional transitions appear in the terahertz (THz) range, where lasers are actively developed to achieve stable frequencies [79], as in the case of quantum cascade lasers [80, 81]. However, the transition to the SAM-LAM coupling mode can reach the wavelengths in the infrared range, where CO2 laser can reach. In the case of the H-containing species, PV-sensitive transitions are observed in the infrared range, such as the transition between #1 and #5 of CHBrIOH presented in Table 1. To the best of our knowledge, these substituted methanol molecules have not been synthesized; however, they may be formulated in the gas phase via laser ablation, which is widely employed (e.g., Refs. 82, 83, 84, 85, 86).
V Conclusions
@@ Acknowledgements.
The author wishes to acknowledge a financial support from the European Union’s Horizon 2022 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101105452 (QDMAP) and the JSPS KAKENHI (grant no 21K14643). The author used the computer resource offered under the category of General Projects by the Research Institute for Information Technology, Kyushu University, and the FUJITSU Supercomputer PRIMEHPC FX1000 and FUJITSU Server PRIMERGY GX2570 (Wisteria/BDEC-01) at the Information Technology Center, The University of Tokyo. The author thanks Edit Mátyus (Budapest) for valuable discussions and Alberto Martń Santa Daría (Salamanca) for the support in adding the PV module to the TENSINT code. The author thanks Editage (www.editage.com) for English language editing.References
@@
- Aucar et al. [2022] I. A. Aucar, Y. Chamorro, and A. Borschevsky, Parity-violating contributions to nuclear spin-rotation interactions and to NMR shielding constants in tetrahedral molecules, Phys. Rev. A 106, 062802 (2022).
- [49] 0.2319 is the default value of the version of the DIRAC code employed in this study. The value of depends on the renormarization prescription, see Ref. 50.
- A Zyla et al. [2020] P. A Zyla, R. M Barnett, J. Beringer, O. Dahl, A. D Dwyer, E. D Groom, C.-J. Lin, K. S Lugovsky, E. Pianori, J. D Robinson, et al., Review of particle physics (rpp 2020), Progress of Theoretical and Experimental Physics 2020, 1 (2020).
@@
- [64] DIRAC, a relativistic ab initio electronic structure program, Release DIRAC22 (2022), written by H. J. Aa. Jensen, R. Bast, A. S. P. Gomes, T. Saue and L. Visscher et al. (available at http://dx.doi.org/10.5281/zenodo.6010450, see also http://www.diracprogram.org), accessed Nov. 5, 2022.
@@
- Thierfelder et al. [2010] C. Thierfelder, G. Rauhut, and P. Schwerdtfeger, Relativistic coupled-cluster study of the parity-violation energy shift of CHFClBr, Phys. Rev. A 81, 032513 (2010).
@@
- Augenbraun et al. [2020] B. L. Augenbraun, J. M. Doyle, T. Zelevinsky, and I. Kozyryev, Molecular asymmetry and optical cycling: Laser cooling asymmetric top molecules, Phys. Rev. X 10, 31022 (2020).
- Persinger et al. [2023] T. D. Persinger, J. Han, A. T. Le, T. C. Steimle, and M. C. Heaven, Electronic spectroscopy and ionization potentials for YbOH and YbOCH3, Phys. Rev. A 107 (2023).
@@
- Becke [1993] A. D. Becke, A new mixing of Hartree–Fock and local density-functional theories, The Journal of chemical physics 98, 1372 (1993).
- Stephens et al. [1994] P. J. Stephens, F. J. Devlin, C. F. Chabalowski, and M. J. Frisch, Ab Initio Calculation of Vibrational Absorption and Circular Dichroism Spectra Using Density Functional Force Fields, J. Phys. Chem. 98, 11623 (1994).
- Yanai et al. [2004] T. Yanai, D. P. Tew, and N. C. Handy, A new hybrid exchange–correlation functional using the coulomb-attenuating method (CAM-B3LYP), Chem. Phys. Lett. 393, 51 (2004).
@@ VI Employed masses
VII Convergence with respect to the DVR grid points
@@ VIII Tight exponents
IX PECs, wavefunctions, and PV curves
@@
Figures 6-10 display the potential energy curves (PECs), 1D torsional wavefunctions, and parity-violating energy () curves of the molecules that are not shown in the main text. In these figures, the blue and green solid lines indicate the odd (ground, 3rd, …) and even-numbered (2nd, 4th, …) torsional state levels. The horizontal axis represents the torsional angle () defined in Eq. 2 of the main text. An example of the direction of the internal rotation of the CXYZ unit is shown in FIG. 3 of the main text. The atomic number increases from atom X to Z, and the wells correspond to , respectively, corresponding to (H-O-C-Y), (H-O-C-Z), and (H-O-C-X), respectively.
Tables 6-10 present the vibrational energies and PV frequencies of the target molecules. In these tables, ( = X, Y, Z = H, F, Cl, Br, I) represents the quantum number that corresponds to the well where the dihedral angle -C-O-H is close to at the corresponding (approximate) values (). “-” refers to mixed states that are delocalized into the three wells.